# Density of wave fronts

### v2

@@ Abstract.
We prove that wave fronts on a flat torus become dense. As a corollary, wave fronts become dense for a square billiard or for the geodesic flow on the flat Klein bottle or the surface of a three dimensional cube.
@@ 1. Motivating questions
Imagine a square room with mirror boundaries. Light emerges from a lamp. The wave front initially forms a perfect circle. After reflecting at the mirror walls, it will bounce back, continuing to be piecewise smooth curve at any time. What happens in the long term with this wave front? Does it fill out the room more and more densely? How long do we have to wait so that every ball of a given fixed radius has and remains to have a non-empty intersection with this curve?
@@
An earth quake has its epicenter at a point on a closed surface. What is the fate of the resulting tsunami wave front emanating from this point? On a round sphere or round projective plane, this wave will stay circular and oscillate forth and back periodically. This situation is obviously very special and already changes if the sphere is changed. What happens for a general surface like an ellipsoid or a torus, surfaces that lack this Wiederkehr property? What happens with a light front radiating from a point on the surface of a cube?
@@
As a third example, imagine a duck in a circular pond, making waves. These waves first form circles around the duck, then bounce off the shore and come back. What happens in the long term? If the duck was in the center of the pond, and assuming ideal conditions, the wave front will reach the shore, reflect there and return back to the duck. Assuming no friction or diffusion, the wave front oscillates forth and back between center and boundary. What happens if the duck is located at a point off the center? Does the wave front produced by the duck become dense in the pond? Intuition says yes because the length of the wave front grows [36]. It gets more complicated as it has to remain wrapped-up in the pond.
@@
These pictures illustrate some rather general wave front problems in Riemannian geometry. One can ask for which Riemannian manifolds with or without boundary or on which polytopes do the wave front become dense for all initial points, like in the mirrored room example? In which cases does the wave front become dense for some points and not become dense for others like in the duck example case? In which cases does it not become dense for all initial points, like in the tsunami case?
@@ 2. Background
The second frame work extends to compact Riemannian manifold $(M,g)$ with boundary $\delta M$. Here, the geodesic flow, referred to as a billiard, requires distinguishing between the motion in the interior of $M$ and reflections at the boundary. The boundary $N=\delta M$ is assumed to be a smooth manifold of co-dimension $1$ with the metric $g$ inducing a Riemannian metric $h$ on $\delta M$. This is achieved by restricting $g$ to the hyper tangent plane $T_{P}N\subset T_{P}M$. If a geodesic $x(t)$ starts at the boundary, it can continue as a boundary geodesic or enter the interior of $M$. A geodesic $x(t)$ starting in the interior hits the boundary at some point $P\in\delta M$, the incoming vector $x^{\prime}(t^{-})\in T_{P}M$ is reflected at the boundary to produce an out-going vector $x^{\prime}(t^{+})$, analogous to a billiard ball rebounding off a table edge or a light ray reflecting off a mirror. In non-convex scenarios, like the Sinai billiard, where a circular obstacle is placed in a flat Clifford 2-torus, the return map to the boundary may already exhibit non-smooth behavior. After each reflection, the geodesic flow resumes in the interior until the next impact point.
@@ 3. Wave fronts
Wave fronts can be defined and observed on any piecewise smooth Riemannian manifold $M$ with or without boundary. The wave front of a point $P$ is defined as the hyper surface $W_{t}=\exp_{P}(S_{t})$, where $\exp_{P}:T_{P}M\to M$ is the exponential map and $S_{t}=\{x\in\mathbb{R}^{d},|x|=t\}$ is the sphere of radius $t$ in the tangent space $T_{P}M\sim\mathbb{R}^{d}$. The global existence of the exponential map on a closed smooth Riemannian manifold is guaranteed by the Hopf-Rynov theorem. In the case of a manifold with boundary, the geodesic flow dynamics is also known as the billiard dynamical system. This is already interesting for a 2-dimensional domain and especially on a convex domains, in which case the system is called the Birkhoff billiard [10, 32]. Already in a polygon like in a square, a path entering one of the corners does not have a well defined bounce but the wave front $W_{t}$ remains defined for all times. Similarly, when looking at a wave front on a surface of a polyhedron like a cube, it is not clear how to continue the geodesic through the corner. We just disregard these trajectories. The wave front $W_{t}$ is still defined now as piecewise smooth curve. It will even become disconnected after finite $t$ in general. In this article, we limit ourselves to
@@
While for a smooth surface $M$ without boundary or a billiard in a convex region, the wave fronts $W_{t}$ are piecewise smooth and in particular continuous, this is already different on a closed polyhedral surface. The wave front there can become disconnected after some time. This happens in particular for the cube surface, the surface bounding the unit cube. Numerical experiments still indicate us also for more general non-flat compact surfaces, that in general, the wave front becomes dense, meaning that it fills out the surface more and more. An unresolved case is already the dodecahedron.
@@
The wave front $W_{t}(P)$ becomes dense on $M$ if for every open ball $U=B_{r}(Q)$ in $M$ with $r>0$, there is a time $\tau(P,r)$ such that for all $t>\tau(P,r)$, the wave front $W_{t}(P)$ intersects $U$ in a non-empty set.
@@
Since the (d1)(d-1)-dimensional volume of the co-dimension-one wave front $W_{t}$ is expected to grow and the manifold $M$ is finite, we expect it to fill up $M$ more and more. The term “generic” is here used in an informal way but it could be made precise using the notion of Baire generic. The set of Riemannian manifolds of a fixed dimension naturally can be made into a metric space by Nash embedding them into a large Euclidean space $E$. A metric on Riemannian manifolds can then be defined by taking the smallest possible Hausdorff distance between any embeddings $M_{1},M_{2}$ in $E$. Alternatively one could fix a topological manifold $M$ and define a distance between Riemannian metrics on $M$ and use this topology on metrics to define Baire genericity.
@@
For a generic $d\geq 2$-dimensional Riemannian manifold with or without boundary, the wave front $W_{t}(P)$ becomes dense for all points $P\in M$.
3.5.
For Wiederkehr manifolds $M$ like the round sphere (or in a generalized sense the projective space), the wave front $W_{t}(P)$ is either a point or a circle for all $t\geq 0$. The story of Wiederkehr manifolds is related to the story of manifolds on which all geodesics are closed [8]. Any round sphere or projective space of constant curvature in any dimension has this property. But this is an exceptional case as already Blaschke realized in 1921 [11]. It is natural to conjecture that for a generic $2$-manifold $M$ with or without boundary, every wave front $W_{t}(P)$ becomes dense in the sense of Definition 1.
@@ 4. Surfaces
We confirm here that one can have dense wave fronts even in very simple situations. Despite the fact that having a dense wave front displays some sort of mixing property. This can even happen in the case when the geodesic flow is integrable. We look first at the flat 2-torus $M=\mathbb{R}^{2}/\mathbb{Z}^{2}$. It is a $2$-dimensional compact Riemannian manifold with Euclidean flat metric, induced from the plane.
@@
Our proof actually shows that for any arc $A=\{(\cos(t),\sin(t)),t\in[\alpha,\beta]$ of positive length, the wave front $W_{t}(P,A)=\exp_{P}(tA)$ becomes dense. We can think of the exponential map restricted to a sector as building a light cone of a headlight. If $\alpha=\beta$, we have a laser, a wave front going in a single direction. In other words this is then a single geodesic ray. While for irrational $\alpha/(2\pi)$ such a laser beam $\{W_{s}(p),s\in[0,t]\}$ becomes dense, the wave front of this laser is just the point $W_{t}(p)$ for all $t$.
Proof.
We look at the wave front on the universal cover $\mathbb{R}^{2}$ of $\mathbb{T}^{2}$, where it is a round circle of radius $t$. We can assume without loss of generality that the point $P$ is at the origin. The wave front can be seen as a union of two half circles which both are graphs. The first is the graph of $f(x)=\sqrt{t^{2}-x^{2}}$. The second is the graph of $g(x)=-\sqrt{t^{2}-x^{2}}$. We forget about the and even on the full graph of $f$, we concentrate only on a small rectangle $R$ of the circle which is chosen to be so wide that its height is $1$. For given $t>0$, define the region $R=[a,b]\times[f(a),f(b)]$ with $a,b]=[-2\cdot\sqrt{t},-\sqrt{2t}]$. The graph of $W_{t}(p)$ intersected with $R$ crosses $R$ monotonically from $(a,f(a))$ to $(b,f(b))$. Since the graph of $f$ is concave down, the slope of $f$ is in the interval $I=[\sqrt{2},2]/\sqrt{1-t/4}$ which is for $t>36/5$ smaller than $3/\sqrt{t}$. We also check using a Taylor expansion of $f$ that that $f(b)-f(a)=1+O(1/\sqrt{t})$. This implies that $|f(t+1)-f(t)|\leq 3/\sqrt{t}$ in that interval $I$. If $R$ is projected to the fundamental domain $0,1]\times[0,1]$, it crisscrosses that square and is $3/\sqrt{t}$ dense. We have shown that even a tiny part $W_{t}(P)\cap R$ of the entire wave front has become $3/\sqrt{t}$ dense. ∎
@@
On a flat Klein bottle $M$, all wave fronts $W_{t}(P)$ become dense.
$4.$
We investigated the torus problem first using results from the Gauss-circle problem which tells how many lattice points we expect in the annulus $A_{t,t+h}$ between $W_{t+h}$ and $W_{t}$ in the universal cover. Already Gauss estimated the number of lattice points in this annulus to be close to its area $2\pi th$ and estimated an error $|E(t)|\leq\sqrt{2}\cdot 2\pi t$. It is known for example that $E(t)\leq Ct^{2/3}$.  [^1] The big open problem in this context is whether $E(t)=O(t^{1/2+\epsilon}$ holds for every $\epsilon>0$. This means that the annulus $A_{t,t+h}$ has about $2\pi th+Ct^{2/3}h$ lattice points. For $h=t^{-1/2}$, this area grows like $2\pi\sqrt{t}$ with an error $O(t^{1/6})$. As we expect so many lattice points in $A_{t,h}$ and every of these lattice points is $h=1/\sqrt{t}$ close to the wave front $W_{t}$ we see that $W_{t}$ gets $1/\sqrt{t}$ close to the original point $P$. Changing the annulus to $A_{t+h,t+2h}$ shows that we have points in distance $1/\sqrt{t},2/\sqrt{t}]$ from $P$. The geometry of numbers approach could even shed better light on how fast the wave fronts fill out the torus. The proof using the Gauss-circle problem in higher dimensions would need heavier mathematical machinery like Ehrhard theory.
@@
The simple constructive proof given above was found only later in the summer of 2024. It illustrates that sometimes, it is easier to prove a stronger result: the proof shows that already a small part of the wave front becomes dense. The reader might enjoy to modify the simple single variable calculus argument to show that also on a stretched torus $\mathbb{R}/(\alpha\mathbb{Z})\times\mathbb{R}/(\beta\mathbb{Z})$, the wave front becomes dense. The argument can also be adapted to wave fronts on higher dimensional flat tori $\mathbb{R}^{d}/\mathbb{T}^{d}$. Again, we just have to restrict to a small cuboid domain $R$, where the wave front can be described in the cover as a graph of a function $f(x_{1},\dots,x_{d-1})$ of d1d-1 variables. There is no doubt that in the case of a non-flat metric, the density of the wave front persists, but this has not been proven; we do not know the shape of the wave front in the universal cover. But it is likely as the wave front gets longer. In the case when the geodesic flow has positive Lyapunov exponents, the wave front length in the universal cover would grow exponentially and make a density result more likely. But we even don’t know how to prove this. In the case when the geodesic flow is hyperbolic, like on a compact Riemannian manifold of strictly negative curvature, the density would follow.
@@ 5. Billiards
There are Wiederkehr situations also for billiards, but it appears to be more restricted. If we start a wave front at the center $P$ of a circular billiard table, then the wave front will periodically return to the point $P$. But already for the same circle, we expect the wave front $W_{t}(P)$ to become dense, if the starting point $P$ is not at the center. We are not aware even whether the density of the wave front in the simple duck problem of the circle has been shown. We can not use the method used here for the torus. In general, we expect:
@@
While the regular equilateral triangle, the regular hexagon or the regular octagon can be looked at using a cover, we do not know what happens for a regular pentagon. It might even hold for non-convex case as long as the illumination problem is solved and the union of wave fronts covers the entire table up to a finite set of points.
@@
We were even unable to find references which ask such wave density questions. But we suspect that it must have been on the mind of many researchers in the field. Some numerical simulations of wave front beams have been done in [6, 5, 7]. These animations show the beauty and complexity which can happen with wave fronts, even for simple situations like a circle or an ellipse. Related to wave fronts are caustics and the “Jacobi last theorem” type question of how many cusps such a caustic has. See [12].

### v4

@@ Abstract.
We prove that wave fronts on a flat torus become dense. As a corollary, wave fronts become dense for a square billiard and for the geodesic flow on the flat Klein bottle and for the surface of a three dimensional cube.
@@ 1. Motivating questions
Imagine a square room with mirror lining the inner walls. Light emerges from a lamp. The wave front initially forms a perfect circle. After reflecting at the mirror walls, it will bounce back, continuing to be piecewise smooth curve at any time. What happens in the long term with this wave front? Does it fill out the room more and more densely? How long do we have to wait so that every ball of a given fixed radius has and remains to have a non-empty intersection with this curve?
@@
An earth quake has its epicenter at a point on a closed surface. What is the fate of the resulting tsunami wave front emanating from this point? On a round sphere or round projective plane, this wave will stay circular and oscillate forth and back periodically. This situation is obviously very special on a perfect sphere. It changes if the metric on the sphere is altered in any way. What happens for a general surface like an ellipsoid or a torus, surfaces that lack this return property? What happens with a light front radiating from a point on the surface of a cube?
@@
As a third example, imagine a duck in a circular pond, making waves. These waves first form circles around the duck, then bounce off the shore and come back. What happens in the long term? If the duck was in the center of the pond, and assuming ideal conditions, the wave front would reach the shore, reflect there and return back to the duck. Assuming no friction or diffusion, the wave front would oscillates back and forth between center and boundary. What would happen if the duck was located at a point off the center point of the pond? Would the wave front produced by the duck become dense on the pond’s surface? Intuition says yes because the length of the wave front grows [37]. It gets more complicated as it has to remain wrapped-up in the pond.
@@
These pictures illustrate some rather general wave front problems in Riemannian geometry. One can ask for which Riemannian manifolds (with or without boundary or on which polytopes) do the wave front become dense for all initial points, like in the mirrored room example? In which cases does the wave front become dense for some points and not become dense for others like in the duck example case? In which cases does it not become dense for all initial points, like in the tsunami case?
@@ 2. Background
The second framework extends to compact Riemannian manifold $(M,g)$ with boundary $\delta M$. Here, the geodesic flow, referred to as a billiard, requires distinguishing between the motion in the interior of $M$ and reflections at the boundary. The boundary $N=\delta M$ is assumed to be a smooth manifold of co-dimension $1$ with the metric $g$ inducing a Riemannian metric $h$ on $\delta M$. This is achieved by restricting $g$ to the hyper tangent plane $T_{P}N\subset T_{P}M$. If a geodesic $x(t)$ starts at the boundary, it can continue as a boundary geodesic or enter the interior of $M$. A geodesic $x(t)$ starting in the interior hits the boundary at some point $P\in\delta M$, the incoming vector $x^{\prime}(t^{-})\in T_{P}M$ is reflected at the boundary to produce an out-going vector $x^{\prime}(t^{+})$, analogous to a billiard ball rebounding off a table edge or a light ray reflecting off a mirror. In non-convex scenarios, like the Sinai billiard, where a circular obstacle is placed in a flat Clifford 2-torus, the return map to the boundary may already exhibit non-smooth behavior. After each reflection, the geodesic flow resumes in the interior until the next impact point. If a geodesics hits a corner, then it is terminated.
@@ 3. Wave fronts
Wave fronts can be defined and observed on any piecewise smooth Riemannian manifold $M$ with or without boundary. The wave front of a point $P$ is defined as the hyper surface $W_{t}=\exp_{P}(S_{t})$, where $\exp_{P}:T_{P}M\to M$ is the exponential map and $S_{t}=\{x\in\mathbb{R}^{d},|x|=t\}$ is the sphere of radius $t$ in the tangent space $T_{P}M\sim\mathbb{R}^{d}$. The global existence of the exponential map on a closed smooth Riemannian manifold is guaranteed by the Hopf-Rynov theorem. In the case of a manifold with boundary, the geodesic flow dynamics is also known as the billiard dynamical system. This is already interesting for a 2-dimensional domain and especially on a convex domains, in which case the system is called the Birkhoff billiard [10, 33]. Already in a polygon, like in a square, a path entering one of the corners does not have a well defined bounce, but the wave front $W_{t}$ remains defined for all times. Similarly, when looking at a wave front on a surface of a polyhedron like a cube, it is not clear how to continue the geodesic through the corner. We just disregard these trajectories. The wave front $W_{t}$ is still defined now as piecewise smooth curve. It will even become disconnected after finite $t$ in general. In this article, we limit ourselves to
@@
While for a smooth surface $M$ without boundary or a billiard in a convex region, the wave fronts $W_{t}$ are piecewise smooth and in particular continuous. This is already different on a closed polyhedral surface. The wave front there can become disconnected after some time. This happens in particular on a cube’s surface, specifically the surface bounding the unit cube. Numerical experiments still indicate to us that also for more general non-flat compact surfaces, in general, the wave front becomes dense, meaning that it fills out the surface more and more. An unresolved case is already the dodecahedron.
@@
We say that the wave front $W_{t}(P)$ becomes dense on $M$ if for every open ball $U=B_{r}(Q)$ in $M$ with $r>0$, there is a time $\tau(P,r)$ such that for all $t>\tau(P,r)$, the wave front $W_{t}(P)$ intersects $U$ in a non-empty set.
@@
To rephrase this, we can also say that a wave front $W_{t}(P)$ becomes dense if for every $\epsilon>0$ there is a time $\tau(P,\epsilon)$ such that for $t>\tau(P,\epsilon)$ the $\epsilon$-neighborhood of $W_{t}(P)$ covers the entire manifold.
3.5.
Since the (d1)(d-1)-dimensional volume of the co-dimension-one wave front $W_{t}$ is expected to grow and the manifold $M$ is finite, we expect it to fill up $M$ more and more. The term “generic” here is used in an informal way, but it could be made precise using the notion of Baire generic. The set of Riemannian manifolds of a fixed dimension naturally can be made into a metric space by Nash embedding them into a large Euclidean space $E$. A metric on Riemannian manifolds can then be defined by taking the smallest possible Hausdorff distance between any embeddings $M_{1},M_{2}$ in $E$. Alternatively one could fix a topological manifold $M$ and define a distance between Riemannian metrics on $M$, using this topology on metrics to define Baire genericity.
@@
For a generic compact $d\geq 2$-dimensional Riemannian manifold with or without boundary, the wave front $W_{t}(P)$ becomes dense for all points $P\in M$.
3.6.
For manifolds $M$ for which all geodesics are closed [8] like the round sphere (or in a generalized sense the projective space), the wave front $W_{t}(P)$ is either a point or a circle for all $t\geq 0$. Any round sphere or projective space of constant curvature in any dimension has this property. But this is an exceptional case as already Blaschke realized in 1921 [11]. It is natural to conjecture that for a generic $2$-manifold $M$ with or without boundary, every wave front $W_{t}(P)$ becomes dense in the sense of Definition 1.
@@ 4. Surfaces
We confirm here that one can have dense wave fronts even in very simple situations, despite the fact that having a dense wave front displays some sort of mixing property. This can happen even in the case when the geodesic flow is integrable. We look first at the flat 2-torus $M=\mathbb{R}^{2}/\mathbb{Z}^{2}$. It is a $2$-dimensional compact Riemannian manifold with Euclidean flat metric, induced from the plane.
@@
We prove a stronger result which states that for a specific arc $A=\{(\cos(\theta),\sin(\theta)),\theta\in[\alpha,\beta]\}$ of positive length, the wave front $W_{t}(P,A)=\exp_{P}(tA)$ becomes dense. Note that the length of $A$ goes to 0. We can think of the exponential map restricted to a sector as building a light cone of a headlight. If $\alpha=\beta$, we have a laser, a wave front going in a single direction. In other words this is then a single geodesic ray. While for irrational $\alpha/(2\pi)$ such a laser beam $\{W_{s}(P),s\in[0,t]\}$ becomes dense, the wave front of this laser is just the point $W_{t}(P)$ for all $t$.
4.3.
In the universal cover $\mathbb{R}^{2}$ of $\mathbb{T}^{2}$, the wave front is a round circle of radius $t$. We can assume without loss of generality that the point $P$ is at the origin. The wave front can be seen as a union of two half circles, both of which are graphs. The first is the graph of $f(x)=\sqrt{t^{2}-x^{2}}$. The second is the graph of $g(x)=-\sqrt{t^{2}-x^{2}}$.
To prove the result, we concentrate only on a small rectangle $R$ containing part of the graph of $f$. It is chosen to be so wide that its height is approximately $1$. For given $t>0$, define the region $R=[a,b]\times[f(a),f(b)]$ with $a,b]=[-2\sqrt{t},-\sqrt{2t}]$. The graph of $W_{t}(P)$ intersected with $R$ crosses $R$ monotonically from $(a,f(a))$ to $(b,f(b))$. Since the graph of $f$ is concave down, the derivative of $f$ is in the interval $\sqrt{2}/\sqrt{t-2},2/\sqrt{t-4}]$ which is for $t>36/5$ smaller than $3/\sqrt{t}$. This implies that $|f(u+1)-f(u)|\leq 3/\sqrt{t}$ for $u\in I$.
If $R$ is projected to $0,1)\times[0,1)$ using $\pi(x,y)=(x\;{\rm mod}1,y\;{\rm mod}1)$, then for every $P=(x,y)\in[0,1)\times[0,1)$, there exists $u\in[a,b]$ such that $\pi(u,\sqrt{t^{2}-u^{2}})-(x,y)||\leq 3/\sqrt{t}$. Proof: Look at all the $u\in[a,b]$ such that $u\;{\rm mod}\;1=x$. As $|f(u+1)-f(u|)\leq 3/\sqrt{t}$, one of the values $u\in\pi^{-1}(x)$ must satisfy $|f(u)-y|\leq 3/\sqrt{t}$ so that $\pi(u,\sqrt{t^{2}-u^{2}})-(x,y)||=\sqrt{0^{2}+|f(u)-y|^{2}}\leq 3/\sqrt{t}$.
We have shown that even a tiny part $W_{t}(P)\cap R$ of the entire wave front has become $3/\sqrt{t}$ dense. So, also the entire wave front is $3/\sqrt{t}$ dense.
@@
On a flat Klein bottle $M$, all the wave fronts $W_{t}(P)$ become dense.
@@
We first investigated the torus problem using results from the Gauss-circle problem,which tells us how many lattice points we expect in the annulus $A_{t,t+h}$ between $W_{t+h}$ and $W_{t}$ in the universal cover. Already, Gauss estimated the number of lattice points in this annulus to be close to its area $2\pi th$ and estimated an error $|E(t)|\leq\sqrt{2}\cdot 2\pi t$. It is known, for example, that $E(t)\leq Ct^{2/3}$.  [^1] The big open problem in this context is whether $E(t)=O(t^{1/2+\epsilon})$ holds for every $\epsilon>0$. This means that the annulus $A_{t,t+h}$ has about $2\pi th+Ct^{2/3}h$ lattice points. For $h=t^{-1/2}$, this area grows like $2\pi\sqrt{t}$ with an error $O(t^{1/6})$. As we expect so many lattice points in $A_{t,h}$ and every of these lattice points is $h=1/\sqrt{t}$ close to the wave front $W_{t}$ we see that $W_{t}$ gets $1/\sqrt{t}$ close to the original point $P$. Changing the annulus to $A_{t+h,t+2h}$ shows that we have points in distance $1/\sqrt{t},2/\sqrt{t}]$ from $P$. The geometry of numbers approach could even shed better light on how fast the wave fronts fill out the torus. The proof using the Gauss-circle problem in higher dimensions would need heavier mathematical machinery like Ehrhard theory.
4.6.
The simple constructive proof given above was found only later in the summer of 2024. It illustrates that sometimes, it is easier to prove a stronger result: the proof shows that already a small part of the wave front becomes dense. The reader might enjoy to modify the simple single variable calculus argument to show that also on a stretched torus $\mathbb{R}/(\alpha\mathbb{Z})\times\mathbb{R}/(\beta\mathbb{Z})$, the wave front becomes dense. The argument can also be adapted to wave fronts on higher dimensional flat tori $\mathbb{R}^{d}/\mathbb{Z}^{d}$. Again, we just have to restrict to a small cuboid domain $R$, where the wave front can be described in the cover as a graph of a function $f(x_{1},\dots,x_{d-1})$ of d1d-1 variables. There is no doubt that in the case of a non-flat metric, the density of the wave front persists. However this has not been proven; we do not know the shape of the wave front in the universal cover. But it is likely as the wave front gets longer. In the case when the geodesic flow has positive Lyapunov exponents, the wave front length in the universal cover would grow exponentially and make a density result more likely. But we even don’t know how to prove this. In the case when the geodesic flow is hyperbolic, like on a compact Riemannian manifold of strictly negative curvature, the density would follow.
@@ 5. Billiards
A pair of points $p,q$ in $M$ is called a Wiedersehen pair, if every geodesic through $p$ passes through $q$ and every geodesic through $q$ passes through $p$. If every point belongs to a Wiedersehen pair, then $M$ is called a “Wiedersehen manifold” and must be a sphere as Blashke has conjectured. Ellipsoids are examples of non-Wiedersehen manifolds, where some Wiedersehen pairs exist. In such situations, not all wave fronts become dense but most are believed to become dense. There are Wiedersehen situations also for billiards. If we start a wave front at the center $P$ of a circular billiard table, then the wave front will periodically return to the point $P$. But already for the same circle, we expect the wave front $W_{t}(P)$ to become dense, if the starting point $P$ is not at the center. We are not aware even whether the density of the wave front in the simple duck problem of the circle has been shown. We can not use the method used here for the torus. In general, we expect:
@@
While the regular equilateral triangle, the regular hexagon the regular octagon or a right angle triangle, can be shown to satisfy the density condition, we do not know what happens already for a regular pentagon. It might even hold for non-convex case as long as the illumination problem is solved there, meaning that union of wave fronts covers the entire table up to a finite set of points. The Penrose table is a non-convex piecewise smooth region with elliptic, circular and straight line boundary parts. No wave front now becomes dense there because placing a light source at any point $P$ leads there to dark regions that can not be reached.
Figure 6. In the Penrose table, there are for every $P$ (at least) two square parts of the table which are not reached by $W_{t}(P)$.
@@
We were unable to find references which even ask such wave density questions. But we suspect that it must have been on the mind of many researchers in the field, as some numerical simulations of wave front beams have been done in [6, 5, 7]. These animations show the beauty and complexity which can happen with wave fronts, even for simple situations like a circle or an ellipse. Related to wave fronts are caustics and the “Jacobi last theorem” type question of how many cusps such a caustic has. See [21].